# Compact superconducting vacuum-gap capacitors with low microwave loss and high mechanical coherence for scalable quantum circuits

### v1

@@ Abstract
Vacuum gap capacitors have recently gained considerable attention in superconducting circuit platforms due to their compact design and low dielectric losses in the microwave regime. Their ability to support mechanical vibrational modes makes them ideal candidates for circuit optomechanics. However, precise control of gap size and achieving high coherence in mechanical modes remain long-standing challenges. Here, we present a detailed fabrication process for scalable vacuum gap capacitors that support ultra-high-coherence mechanical motion, exhibit low microwave loss, and maintain a small footprint compared to planar geometries. We fabricate arrays of up to 24 LC resonators, with capacitors featuring nanometer-scale gap size variations. We demonstrate that the mechanical quality factors can reach up to $40\times 10^{6}$, a 100-fold improvement over other platforms, with microwave quality factors $\mathcal{O}(10^{5})$ at low photon number levels. This platform also achieves a sizable single-photon optomechanical coupling rate of approximately 20Hz. Using this, we cooled the mechanical oscillator to its ground state (0.07 quanta) and squeezed its motion below the vacuum level by 2.7dB [1]. We further demonstrate the scalability of this platform by implementing large-scale optomechanical arrays, a strained graphene model [2], and observing quantum collective phenomena in a mechanical hexamer [3]. These vacuum gap capacitors are promising candidates for coupling superconducting qubits with mechanical systems, serving as storage elements in quantum computing, and exploring gravitational effects on quantum mechanics.
@@ 1.   Introduction
In this work, we present a comprehensive account of the detailed fabrication process for low-loss vacuum gap capacitors, tailored for circuit quantum optomechanics and circuit quantum electrodynamics applications, and outline the challenges we encountered. We introduce the ”flat-geometry” vacuum gap capacitor, developed to address the limitations of the conventional platform [37, 38, 29, 39], and describe the full optimization process. As shown in Fig. 1a, the flat geometry offers significant advantages, enabling precise control of both the gap and the mechanical frequency through the lithographic process. In Fig. 1b, we provide a SEM micrograph of the top electrode of the vacuum gap capacitor. Furthermore, Fig. 1c illustrates the performance of our fabrication method compared to other approaches reported in the literature. Finally, we demonstrate the application of this device in a circuit optomechanical platform, achieving accurate target frequencies and quality factors up to 40 million. This platform has facilitated the realization of topological lattices in optomechanical circuits [2], the preparation of squeezed mechanical states and the observation of their decoherence in real time [1], and, most recently, collective ground-state cooling [3].
@@ 2.   Flat geometry vacuum gap capacitors
The conventional fabrication process of vacuum gap relies on deposition and following lithography definition of a sacrificial layer covering the bottom layer. Following this step, several sacrificial materials on different substrates such as Si3N4 on sapphire [37], polymer on Si [38], SiO2 on quartz [29], and a-Si on sapphire [39] have been tested. In these platforms, due to the deposition-induced compressive stress in the superconducting thin film, the drumhead capacitor buckles up after the release at room temperature, which increases the gap size between two plates up to a few micrometers. Cooling down such devices induces tensile stress in the thin film metal due to the significant difference in thermal expansion rates between the thin film and the substrate. Under tensile stress, the drumhead shrinks, resulting in an approximately 50 nm gap size that is neither predictable nor reproducible. This prevents the precise control of the microwave and mechanical properties of the system at low temperatures and reduces the reproducibility of the design given the high probability of deformations and collapses after the release [40]. In practice, any non-uniformity of the stress distribution in the drumhead after release or asymmetric buckling results in an uncertainty in the final gap size at cryogenic temperatures. This can be in the order of tens of nano-meters, hence limiting the frequency fluctuation in the microwave LC resonator in the order of $\mathcal{O}(10\%)$.
@@
High coherence superconducting circuits are traditionally realized in aluminum [42], niobium [43, 44] and more recently tantalum [45, 46]. With the aim of integrating vacuum gap capacitor in optomechanical system, we decide to use aluminum due to its lighter density which provides larger optomechanical coupling (see Sec. 5 for more details). This makes future integration of our circuits with conventional superconducting qubits more straightforward.
For the sacrificial layer, amorphous silicon (a-Si), silicon nitride (Si3N4), silicon oxide (SiO2), and polymer photoresists are four candidates which were used in the previous generation of circuit optomechanical devices. Each one needs different isotropic etching for the release process. For example, a-Si can be removed by XeF2, which is an exothermic gas etching. Si3N4 can be etched by SF6 plasma, and polymer resists by oxygen plasma. We decided to use SiO2 as our sacrificial layer which can be removed with HF vapor, enabling us to release high aspect ratio structures by avoiding plasma or wet etching, therefore increasing the yield and successful release rate of the process. More details will be discussed in the release section (Sec. 3.5).
@@ 3.1 Etching trenches in silicon
We use optical lithography to transfer patterns on the photoresist. It is performed by direct mask-less optical lithography (${}^{\text{\textregistered}}$ MLA 150). We spin coat a 1 $\mu$m thick ${}^{\text{\textregistered}}$ ECI 3007 positive photoresist after HMDS surface preparation. All photoresist coating and developing steps are processed using automatic coater/developer (${}^{\text{\textregistered}}$ ACS200 GEN3). The exposure dose and depth of focus vary based on the tool and need to be calibrated by dose tests, but are typically set to $\sim~{}150$ mJ/cm2. After the exposure, the resist is developed, and the wafer is rinsed in a spin dryer to clean any unwanted contamination. To remove residual photoresist on the surface of exposed areas, we conduct a short (10-20 seconds) oxygen plasma descum at 200 Watts and 200 sccm (${}^{\text{\textregistered}}$ GiGAbatch). After the descum, the wafer is ready for the etching step.
@@
After each etching step, the photoresist is stripped first using UFT remover 1165 wet process, followed by rinse and drying, and then 3 minutes 600 Watt and 400 sccm Oxygen plasma (${}^{\text{\textregistered}}$ GiGAbatch). In the fabrication steps where the wafer contains uncovered thin-film aluminum, it is recommended to reduce either the power or exposure time of the oxygen plasma to avoid additional oxidation and local heating of the metal, specifically for the vibrating top plate.
@@ 3.2 Bottom layer
The first electrode of the vacuum gap capacitor is placed within the silicon trenches. We also add a spiral inductor that will form an LC resonator. For plane circuits, the metal-substrate interface has a major effect on the total loss of the superconducting circuit [47]. While the electric field in the vacuum gap capacitor is mainly stored in the space between the two electrodes, we still clean the wafer after trench etching with Piranha and dip it into HF (1% diluted) for a few minutes to remove the native silicon oxide and minimize as much as possible dielectric loss. Then we rinse and dry the wafer and immediately transfer it to the deposition tool (less than 3 minutes) and pump down the chamber to avoid regrowth of the native oxide.
Deposition of the bottom aluminum layer can be done by either sputtering or electron beam evaporation. However, we find that evaporated films have better thickness uniformity and thickness control compared to sputtered ones. We typically choose 100 nm thickness (Alliance-Concept EVA 760) for the bottom layer.
@@ 3.3 Sacrificial layer
Figure 4: SiO2 sacrificial layer deposition. a and b, False color SEM micrograph of the cross section of low temperature oxide low pressure chemical vapor deposition (LTO LPCVD) of SiO2 sacrificial layer covering the trenches. The step coverage and gap filling is perfect, the porosity is low, and the oxide layer is dense. c and d, Plasma enhanced chemical vapor deposition (PECVD) of oxide at 200 Celsius. The oxide layer is porous and does not show a good step coverage, forming void areas at the corners of the trench.
@@
We find that there is a trade-off between uniformity and residual topography after polishing: longer times reduce the topography (that finally saturates by dishing effect) but increases the thickness non-uniformity at wafer scale. In order to minimize dishing and delamination effect and increase the polishing uniformity, we fill all the empty areas of the wafer between circuits with dummy patterns. These patterns are squares with the size of $60\mu$m with double of this size spacing (see SI for mode details). In Fig. 5, we plot the mechanical profilometer (${}^{\text{\textregistered}}$ Tencor D600) traces across trenches after each of these cycles (with arbitrary translation to make the different profiles comparable). After 10 minutes, all topography is removed and the roughness is smaller than 10 nm as shown in the inset (see SI for more details)
@@ 3.4 Top aluminum layer deposition
To overcome this drawback, we argon-ion mill and deposit the metal without breaking the vacuum in Plassys. This can help to significantly reduce the microwave cavity heating when sending high power to the device [1]. An optimized evaporation on 200 nm Al in Plassys gives minor tensile stress in the film. Finally, an Al2O3 layer is grown on the drumhead resonator by injecting 10 mBar of 99.99% pure oxygen in Plassys oxidation chamber. Additional deposition parameters are discussed in Sec. 5.
@@ 3.5 Release
We dice the wafer into chips before the release (see details on dicing in SI). The last step of the fabrication process is releasing the vacuum gap capacitors by removing all $\rm{SiO_{2}}$ sacrificial layer from trenches. This is done by vapor phase Hydrofluoric acid etching (${}^{\text{\textregistered}}$ uEtch). Reaction with the sacrificial SiO2 on the wafer surface produces silicon tetrafluoride (SiF4) gas and water vapor:
@@
Although the liquid HF attacks aluminum, we observe that the vapor HF does not deteriorate or degrade Al films. We use a recipe with 125 Torr pressure and $\sim 100$ nm/minute etch rate for 900 seconds etching time in every cycle. Since the vapor HF needs to penetrate horizontally between the top Al layer and the trench bottom surface, the total number of cycles needed should be calculated based on the maximum lateral distance between two penetration windows for the gas etchant, considering the pattern of Al covering the trenches. We multiply this number by a factor of four to prevent any residual oxide and to ensure the whole structure is released. The etching process is liquid-free, which is crucial to release structures with high aspect ratios, in our case $\sim 100$ $\mu$m big drums suspended by $\sim 200$ nm gap above another metallic layer. Any water formation will result in the collapse and sticking of two capacitor plates.
Examples of successfully released devices are shown in Figs. 9 and 10.
Figure 10: Elements of the final device. a, Focused ion beam cross-section of a capacitor before removing the SiO2 sacrificial layer (Pt is used as the focused ion beam protective layer). The flatness of the top layer is visible in the image indicating a successful CMP planarization. b, SEM image of the perforated released drumhead. The holes facilitate the release process. c, Magnified SEM of one release hole. The bottom Al layer is visible from the hole. The small particles seen on the aluminum surface are aluminum hillocks, a well-know accumulation of aluminum in evaporation technique. The size and distribution of hillocks depends on the evaporation rate and the pressure of the chamber. d, The bottom wire going under a drum. e, A magnified SEM of a drumhead clamp. f and g, The spiral inductor air bridges and the galvanic connection.
To facilitate the release, specifically for big drums, we perforate drums by defining small holes with 1.8 $\mu$m diameter and distance of $\sim 10\mu$m as shown in Fig. S14.
We successfully release vacuum gaps with the gap size down to 75 nm, but the success rate of the release was not high for such a low gap size (Fig. S13). For gaps above 150 nm, we find almost 100% yield of release. The success rate depends on the trench size, the thickness of the top plate, and the room temperature stress. Smaller drums have a lower risk of collapse and can tolerate smaller gap sizes. We usually use 150-200 nm thick Al top layer. Attempts to release thin top layers (50 nm) were not successful (see Fig. S15).
@@ 4.   Measurement and results
The bare mechanical damping rates can be either characterized by extracting the linewidth of the OMIT features while pumping the microwave resonance with low powers corresponding to optomechanical cooperativities below one ($\mathcal{C}\ll 1$) or conducting a ring-down experiment. Since the presented fabrication process results in sub-Hertz damping rates of mechanical modes, the OMIT measurement requires low frequency resolution and typically slow sweeping speed. Therefore, the ringdown measurement is more suitable for such characterizations. For higher red-detuned probe powers the effective mechanical damping rate is $\Gamma_{\mathrm{tot}}=\Gamma_{\mathrm{m}}(1+\mathcal{C})$. Sweeping the power of the red probe enables us to directly measure $\Gamma_{\mathrm{m}}$ (Figs. 11c and d).
@@
A persistent challenge in circuit optomechanical platforms has been the heating of the microwave cavity caused by high circulating currents in the circuit [10], which limits the minimum occupation of the mechanical oscillator using sideband cooling. As discussed in Sec. 3.4, argon milling of the bottom-layer aluminum prior to making the galvanic connection removes the resistive aluminum oxide layer, thereby reducing microwave loss. Figure 12 shows the microwave cavity heating with and without ion milling treatment, demonstrating that this method reduces heating by an order of magnitude. With this improvement, the microwave resonator quality factor is limited by dielectric loss of its capacitive element. By measuring the internal microwave quality factor as a function of the intracavity photon number, we report internal quality factors up to $10^{5}$ for photon number $n\sim\mathcal{O}(100)$ going up to $2\times 10^{6}$ for photon number $n\sim\mathcal{O}(10^{7})$ for a resonance frequency $\sim\mathcal{O}(5\,\rm{GHz})$.
@@
Optomechanical building blocks can be couple with each other to form arrays and lattices [52, 53, 54, 55, 56, 57]. Namely, such multimode systems can implement Su-Schrieffer-Heeger (SSH) arrays [58, 59], as a fundamental topological model. Such optomechanical arrays can exhibit non-trivial topological properties, localized edge states, and be exploited to directly measure hybridized microwave modeshapes to fully reconstruct the system’s Hamiltonian [2]. To demonstrate the scalability of our process, we fabricated a 12-site 1D array of identical electromechanical LC circuits with alternating mutual inductance realizing SSH model (Fig. 15a). We slightly swept the trench radius of drumhead resonators in the array (by 500 nm) to shift their frequencies and be able to identify them in the OMIT response (Fig. 15b). As shown in Fig. 15c, the measured mechanical frequencies from OMIT experiment perfectly follow the theoretical fit. The standard deviation of mechanical frequency disorder from the theoretical fit is less than 0.2%, which shows a perfect control and high reproducibility in the process.

### v2

@@ Abstract
Vacuum-gap capacitors have recently attracted significant interest in superconducting circuit platforms due to their compact design and exceptionally low dielectric losses in the microwave regime. Their intrinsic ability to support mechanical vibrational modes makes them well-suited for circuit optomechanics. However, precise control over the gap size and the realization of high-coherence mechanical modes remain longstanding challenges. Here, we present a detailed and scalable fabrication process for vacuum-gap capacitors that support ultra-high-coherence mechanical motion, exhibit low microwave loss, and occupy a significantly smaller footprint compared to conventional planar geometries. By employing a planarized SiO2 sacrificial layer, we achieve vacuum gaps on the order of 150 nm. Using this platform, we have recently demonstrated ground-state cooling and motion squeezing of a mechanical oscillator with a quality factor of 40 million—a 100-fold improvement compared to prior works—as well as a single-photon optomechanical coupling rate of approximately 15Hz [1]. Additional achievements include the realization of an optomechanical topological lattice with 24 sites [2] and the observation of quantum collective dynamics in a mechanical hexamer [3]. Collectively, these results underscore the potential of vacuum-gap capacitors as a platform for coupling superconducting qubits to mechanical systems, enabling quantum storage, and probing gravitational effects in quantum mechanics.
@@ 1.   Introduction
In this work, we present a comprehensive account of the detailed fabrication process for low-loss vacuum gap capacitors, tailored for circuit quantum optomechanics and circuit quantum electrodynamics applications, and outline the challenges we encountered. We introduce the “flat-geometry” vacuum gap capacitor, developed to address the limitations of the conventional platform [37, 38, 29, 39], and describe the full optimization process. As shown in Fig. 1a, the flat geometry offers significant advantages, enabling precise control of both the gap and the mechanical frequency through the lithographic process. In Fig. 1b, we provide a SEM micrograph of the top electrode of the vacuum gap capacitor. Furthermore, Fig. 1c illustrates the performance of our fabrication method compared to other approaches reported in the literature, excluding metalized membranes that feature very large footprint [40, 41] (see supplementary information). Finally, we demonstrate the application of this device in a circuit optomechanical platform, achieving accurate target frequencies and quality factors as high as 40 million. This platform has facilitated the realization of topological lattices in optomechanical circuits [2], the preparation of squeezed mechanical states and the observation of their decoherence in real time [1], and, most recently, collective ground-state cooling [3].
@@ 2.   Flat geometry vacuum gap capacitors
The conventional fabrication process of vacuum gap relies on deposition and following lithography definition of a sacrificial layer covering the bottom layer. Several sacrificial materials on different substrates such as Si3N4 on sapphire [37], polymer on Si [38], SiO2 on quartz [29], and a-Si on sapphire [39] have been tested. In these platforms, due to the deposition-induced compressive stress in the superconducting thin film, the drumhead capacitor buckles up after the release at room temperature, which increases the gap size between two plates up to a few micrometers. Cooling down such devices induces tensile stress in the thin film metal due to the significant difference in thermal expansion rates between the thin film and the substrate. Under tensile stress, the drumhead shrinks, resulting in an approximately 50 nm gap size that is neither predictable nor reproducible. This prevents the precise control of the microwave and mechanical properties of the system at low temperatures and reduces the reproducibility of the design given the high probability of deformations and collapses after the release [42]. In practice, any non-uniformity of the stress distribution in the drumhead after release or asymmetric buckling results in an uncertainty in the final gap size at cryogenic temperatures. This can be in the order of tens of nano-meters, hence limiting the frequency fluctuation in the microwave LC resonator in the order of $\mathcal{O}(10\%)$.
@@
High coherence superconducting circuits are traditionally realized in aluminum [44], niobium [45, 46] and more recently tantalum [47, 48]. With the aim of integrating vacuum gap capacitor in optomechanical system, we decide to use aluminum due to its lighter density which provides larger optomechanical coupling (see SI for more details). This makes future integration of our circuits with conventional superconducting qubits more straightforward.
For the sacrificial layer, amorphous silicon (a-Si), silicon nitride (Si3N4), silicon oxide (SiO2), and polymer photoresists are four candidates which were used in the previous generation of circuit optomechanical devices. Each one needs different isotropic etching for the release process. For example, a-Si can be removed by XeF2, which is an exothermic gas etching. Si3N4 can be etched by SF6 plasma, and polymer resists by oxygen plasma. We decided to use SiO2 as our sacrificial layer which can be removed with HF vapor, enabling us to release high aspect ratio structures by avoiding plasma or wet etching, therefore increasing the yield and successful release rate of the process. In addition, it has infinite selectivity to aluminum and silicon. More details are given in the release section (Sec. 3.5).
@@ 3.1 Etching trenches in silicon
We use optical lithography to transfer patterns on the photoresist. It is performed by direct mask-less optical lithography (${}^{\text{\textregistered}}$ MLA 150). We spin coat a 1 $\mu$m thick ${}^{\text{\textregistered}}$ ECI 3007 positive photoresist after HMDS surface preparation. All photoresist coating and developing steps are processed using automatic coater/developer (${}^{\text{\textregistered}}$ ACS200 GEN3). The exposure dose and depth of focus vary based on the tool and need to be calibrated by dose tests, but are typically set to $\sim~{}150$ mJ/cm2 and 0, respectively. After the exposure, the resist is developed, and the wafer is rinsed in a spin dryer to clean any unwanted contamination. To remove residual photoresist on the surface of exposed areas, we conduct a short (10-20 seconds) oxygen plasma descum at 200 Watts and 200 sccm (${}^{\text{\textregistered}}$ GiGAbatch). After the descum, the wafer is ready for the etching step.
@@
After each etching step, the photoresist is stripped first using UFT remover 1165 wet process, followed by rinse and drying, and then 3 minutes 200 Watt and 200 sccm Oxygen plasma (${}^{\text{\textregistered}}$ GiGAbatch). In the fabrication steps where the wafer contains uncovered thin-film aluminum, it is recommended to reduce either the power or exposure time of the oxygen plasma to avoid additional oxidation and local heating of the metal, specifically for the vibrating top plate.
@@ 3.2 Bottom layer
The first electrode of the vacuum gap capacitor is placed within the silicon trenches. We also add a spiral inductor that will form an LC resonator. For microwave circuits, the metal-substrate interface has a major effect on the total loss of the superconducting circuit [49]. While the electric field in the vacuum gap capacitor is mainly stored in the space between the two electrodes, we still clean the wafer after trench etching with Piranha and dip it into HF (1% diluted) for a few minutes to remove the native silicon oxide and minimize as much as possible dielectric loss. Then we rinse and dry the wafer and immediately transfer it to the deposition tool (less than 3 minutes) and pump down the chamber to avoid regrowth of the native oxide.
Deposition of the bottom aluminum layer can be done by either sputtering or electron beam evaporation. However, we find that evaporated films have better thickness uniformity and thickness control compared to sputtered ones. We typically choose 100 nm thickness (Alliance-Concept EVA 760) for the bottom layer, deposited with 0.5 nm/sec rate.
@@ 3.3 Sacrificial layer
Figure 4: SiO2 sacrificial layer deposition. a and b, False color SEM micrograph of the cross section of low temperature oxide low pressure chemical vapor deposition (LTO LPCVD) of SiO2 sacrificial layer covering the trenches. The step coverage and gap filling is perfect, the porosity is low, and the oxide layer is dense. c and d, Plasma enhanced chemical vapor deposition (PECVD) of oxide at ${200^{\circ}}{\rm{C}}$. The oxide layer is porous and does not show a good step coverage, forming void areas at the corners of the trench.
@@
We find that there is a trade-off between uniformity and residual topography after polishing: longer times reduce the topography (that finally saturates by dishing effect) but increases the thickness non-uniformity at wafer scale. In order to minimize dishing and delamination effect and increase the polishing uniformity, we fill all the empty areas of the wafer between circuits with dummy patterns. These patterns are squares with the size of $60\mu$m with double of this size spacing (see SI for more details). In Fig. 5, we plot the mechanical profilometer (${}^{\text{\textregistered}}$ Tencor D600) traces across trenches after each of these cycles (with arbitrary translation to make the different profiles comparable). After 10 minutes, topography is removed to below 10 nm, as it is shown in the inset (see SI for more details).
@@ 3.4 Top aluminum layer deposition
To overcome this drawback, we argon-ion mill and deposit the metal without breaking the vacuum in Plassys. This can help to significantly reduce the microwave cavity heating when sending high power to the device [1]. An optimized evaporation on 200 nm Al in Plassys gives minor tensile stress in the film. Finally, an Al2O3 layer is grown on the drumhead resonator by injecting 10 mBar of 99.99% pure oxygen in Plassys oxidation chamber. Additional deposition parameters are discussed in supplementary information (SI).
@@ 3.5 Release
We dice the wafer into chips before the release (see details on dicing in SI). The last step of the fabrication process is releasing the vacuum gap capacitors by removing the $\rm{SiO_{2}}$ sacrificial layer. This is done by vapor phase Hydrofluoric acid etching (${}^{\text{\textregistered}}$ uEtch). Reaction with the sacrificial SiO2 on the wafer surface (in the presence of ethanol as the catalyst) produces silicon tetrafluoride (SiF4) gas and water vapor:
@@
Although the liquid HF attacks aluminum, we observe that the vapor HF does not deteriorate or degrade Al films. We use a recipe with 125 Torr pressure and $\sim 100$ nm/minute etch rate for 900 seconds etching time in every cycle. Since the vapor HF needs to penetrate horizontally between the top Al layer and the trench bottom surface, the total number of cycles needed should be calculated based on the maximum lateral distance between two penetration windows for the gas etchant, considering the pattern of Al covering the trenches. We multiply this number by a factor of four to prevent any residual oxide and to ensure the whole structure is released. The etching process is liquid-free, which is crucial to release structures with high aspect ratios, in our case $\sim 100$ $\mu$m big drums suspended by $\sim 200$ nm gap above another metallic layer. Any liquid formation will result in the collapse and sticking of two capacitor plates. Examples of successfully released devices are shown in Figs. 9 and 10.
Figure 10: Elements of the final device. a, SEM micrograph of a focused ion beam cross-section of a capacitor before removing the SiO2 sacrificial layer (Pt is used as the focused ion beam protective layer). The flatness of the top layer is visible in the image indicating a successful CMP planarization. b, SEM image of the perforated released drumhead. The holes facilitate the release process. c, Magnified SEM of one release hole. The bottom Al layer is visible from the hole. The small particles seen on the aluminum surface are aluminum hillocks, a well-know accumulation of aluminum in evaporation technique. The size and distribution of hillocks depends on the evaporation rate and the pressure of the chamber. d, The bottom wire going under a drum. e, A magnified SEM of a drumhead clamp. f and g, The spiral inductor air bridges and the galvanic connection.
To facilitate the release, specifically for big drums, we perforate drums by defining small holes with 1.8 $\mu$m diameter and distance of $\sim 10\mu$m as shown in Fig. 10. We successfully release vacuum gaps with the gap size down to 75 nm, but the success rate of the release was not high for such a low gap size (see SI). For gaps above 150 nm, we find almost 100% yield of release. The success rate depends on the trench size, the thickness of the top plate, and the room temperature stress. Smaller drums have a lower risk of collapse and can tolerate smaller gap sizes. We usually use 150-200 nm thick Al top layer. Attempts to release thin top layers (50 nm) were not successful (see SI).
@@ 4.   Measurement and results
The bare mechanical damping rates can be either characterized by extracting the linewidth of the OMIT features while pumping the microwave resonance with low powers corresponding to optomechanical cooperativities below one ($\mathcal{C}\ll 1$) or conducting a ring-down experiment. Since the presented fabrication process results in sub-Hertz damping rates of mechanical modes, the OMIT measurement requires low frequency resolution and typically low resolution bandwidth. Therefore, the ringdown measurement is more suitable for such characterizations. For higher red-detuned probe powers, the effective mechanical damping rate is $\Gamma_{\mathrm{tot}}=\Gamma_{\mathrm{m}}(1+\mathcal{C})$. Sweeping the power of the red probe enables us to directly measure $\Gamma_{\mathrm{m}}$ (Figs. 11c and d).
@@
In Fig. 13a, we reported the mechanical quality factors of 14 devices fabricated on a single chip. In Fig. 14, we present data from a much broader dataset comprising 119 mechanical oscillators measured over a span of three years, encompassing three different fabrication batches across six wafers. The fabrication of batches CCv3, CCv4, and CSv1 was completed in September 2022, May 2023, and September 2024, respectively. The average mechanical quality factor across 119 devices is remarkably high, at 4.3 million. These results demonstrate the high yield and consistency of our fabrication process across batches and over time.
A persistent challenge in circuit optomechanical platforms has been the heating of the microwave cavity caused mainly by high circulating currents in the circuit [10], which limits the minimum occupation of the mechanical oscillator using sideband cooling. As discussed in Sec. 3.4, argon milling of the bottom-layer aluminum prior to making the galvanic connection removes the resistive aluminum oxide layer, thereby reducing microwave loss. Figure 12 shows the microwave cavity heating with and without ion milling treatment, demonstrating that this method reduces heating by an order of magnitude. With this improvement, the microwave resonator quality factor is limited by dielectric loss of its capacitive element. By measuring the internal microwave quality factor as a function of the intracavity photon number, in [3] we report internal quality factors up to $10^{5}$ for photon number $n\sim\mathcal{O}(100)$ going up to $2\times 10^{6}$ for photon number $n\sim\mathcal{O}(10^{7})$ for a resonance frequency $\sim\mathcal{O}(5\,\rm{GHz})$. The reported microwave quality factor is compatible with the large participation ratio of the electric field in the native aluminum oxide present on the capacitor leads and other vacuum gap capacitors have shown similar performance in compact capacitor for resonator [52] and for transmon qubit [53].
@@
Figure 14: Batch-to-batch mechanical quality factors. The mechanical quality factors for three different batches CCv3, CCv4, and CSv1, which are fabricated on September 2022, May 2023, and September 2024, respectively. The transparent region is the estimated density of quality factors, the top line show the maximum quality factor, the middle line shows the average quality factor, and the bottom line shows the minimum quality factor. All the drums have 70 $\mu$m radius (average mechanical frequency 2 MHz).
@@
Optomechanical building blocks can be coupled with each other to form arrays and lattices [56, 57, 58, 59, 60, 61]. Namely, such multimode systems can implement Su-Schrieffer-Heeger (SSH) arrays [62, 63], as a fundamental topological model. Such optomechanical arrays can exhibit non-trivial topological properties, localized edge states, and be exploited to directly measure hybridized microwave modeshapes to fully reconstruct the system’s Hamiltonian [2]. To demonstrate the scalability of our process, we fabricated a 12-site 1D array of identical electromechanical LC circuits with alternating mutual inductance realizing SSH model (Fig. 16a). We slightly sweep the trench radius of drumhead resonators in the array (by 500 nm) to shift their frequencies and be able to identify them in the OMIT response (Fig. 16b). As shown in Fig. 16c, the measured mechanical frequencies from OMIT experiment perfectly follow the theoretical fit. The standard deviation of mechanical frequency disorder from the theoretical fit is less than 0.2%, which shows a perfect control and high reproducibility in the process.